OpenStudy (anonymous):

Solve 3(x-2)+6=4(2-x) .

OpenStudy (anonymous):

So it's all understood then, tierram?

OpenStudy (anonymous):

We hope! ha

OpenStudy (anonymous):

I am getting \( x = \frac 8 7 \)

OpenStudy (anonymous):

Will, you solution is incorrect.

OpenStudy (anonymous):

aye. god i should not stay up all night.

OpenStudy (anonymous):

Multiply out the brackets on each side. 3x - 6 + 6 = 8 - 4x Simply each side. 3x = 8 - 4x Move x's to one side. 7x = 8 => x = (8/7)

OpenStudy (anonymous):

\[\Large \begin{array}{l} 3(x - 2) + 6 = 4(2 - x)\\ 3x - 6 + 6 = 8 - 4x\\ 7x = 8\\ x = \frac{8}{7} \end{array}\] Yeah. FFM is right. I missed the -4x the first time.